A gaseous mixture consists of 1.27 mol of helium, 0.36 mol of neon, and 1.58 mol of argon.
What is the mole fraction of helium in the mixture?
1
Expert's answer
2019-12-09T08:05:52-0500

æ - mole fraction, n -number of molesæcomponents=ncomponents / "\sum" nall components "\sum" æcomponents=1æHe=nHe/nNe+nHe +nAræHe=0,395